112 mL of NO2(g) at STP was liquefied, the density of the liquid being 1.15 g mL
Calculate the volume of liquid and the number of molecules in the liquid NO2
1
Expert's answer
2020-08-05T05:06:26-0400

One mole of an ideal gas will occupy a volume of 22.4 liters at STP (Standard Temperature and Pressure, 0°C and one atmosphere pressure).

Proportion:

1 mole – 22400 ml

n moles – 112 ml

n = 0.005 moles of NO2

m=n×M

M(NO2) = 46 g/mol

m(NO2) = 0.005×46=0.23 g

ρ(NO2) =1.15 g/ml

V = m/ρ

V = 0.23/1.15 = 0.2 ml

One mole of NO2 molecules contains 6.02×1023 NO2 molecules

Number of molecules = n×6.02×1023 = 0.005×6.02×1023 = 3.01×1021 molecules

Answer: 0.2 ml and 3.01×1021 molecules